The standard enthalpy change of combustion [to CO2(g) and H2O()] at 25°C of the organic solid bianthracene, C28H18(s), is determined to be -13924.1 kJ mol-1. What is the Hf° of C28H18(s) based on this value?

Use the following data: Hf° H2O () = -285.83 kJ mol-1 ; Hf° CO2(g) = -393.51 kJ mol-1


_________ kJ mol-1
Answer #1

C28H18(s) + 67/2 O2\rightarrow 28CO2 + 9 H2O. ∆H = -13924.1kJ/mol

∆H = enthalpy of formation of products - enthalpy of formation of reactants

-13924.1 = 28× (-393.51) + 9×(-285.83) - ( 67/2 × 0 + ∆H°f,bianthracene )

∆H°f = -13590.75 + 13924.1 = 333.35kJ/mol

= 333.35 kJ/mol. (Answer)